OpenStudy (anonymous):

What's an equation of the line with a slope of -5 and a y-intercept of (0,3)

OpenStudy (anonymous):

the "slope intercept" form is \[y=mx+b\] in your case \(m=-5,b=3\)

OpenStudy (anonymous):

put them directly in to the formula to get your answer okay with this?

OpenStudy (anonymous):

What does x represent in this equation?

OpenStudy (anonymous):

\(x\) is a variable there is an \(x\) in your answer

OpenStudy (anonymous):

just like before when you had to find the range of \(y=3x-2\) for different values of \(x\) the \(x\) can be any number

OpenStudy (anonymous):

in this case if you put \(m=-5\) and \(b=3\) you get a similar equation \[y=-5x+3\]

OpenStudy (anonymous):

so for example if \(x=2\) you would get \(y=-5\times 2+3=-10+3=-7\) telling you that te point \((2,-7)\) is on the graph

OpenStudy (anonymous):

@satellite73 What if my problem is like this.... Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

then you have to find the slope first

OpenStudy (anonymous):

\[m=\frac{-4-2}{0-(-2)}=\frac{-6}{2}=-3\]

OpenStudy (anonymous):

now you already have the \(y\) intercept is \(-4\) because you are told \((0,-4)\) is on the graph

OpenStudy (anonymous):

use \[y=mx+b\] with \(m=-3,b=-4\)

OpenStudy (anonymous):

y=-3x-4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think I got the hang of this... once again thank you so much @satellite73. You've been the best!

OpenStudy (anonymous):

you are quite welcome